Required: Calculate Accumulated Depreciation - Machine at December 31, 2018 Using Declining Balance                          
                          
                          
Data:                          
Cost of machine   $800,000                       
Estimated useful life   10 Years                      
Estimated salvage value   $150,000                       
Double Declining   200%                      
Purchase Date    January 1, 2017                      
Answer #1

Answer

  • Requirement: Accumulated Depreciation – Machine at Dec 31, 2018, which is the end of SECOND YEAR from purchase
  • Working

A

Cost

$          800,000.00

B

Residual Value

$          150,000.00

C=A - B

Depreciable base

$          650,000.00

D

Life [in years]

10

E=C/D

Annual SLM depreciation

$            65,000.00

F=E/C

SLM Rate

10.00%

G=F x 2

DDB Rate

20.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

Accumulated Depreciation

2017

$           800,000.00

20.00%

$       160,000.00

$           640,000.00

$      160,000.00

2018

$           640,000.00

20.00%

$       128,000.00

$           512,000.00

$      288,000.00 = Answer

  • Answer:
    Accumulated Depreciation – Machine at Dec 31, 2018 = $ 288,000